OpenStudy (anonymous):

Which is the 25th term of the arithmetic sequence where a1 = 8 and a9 = 80 ?

OpenStudy (anonymous):

224

OpenStudy (anonymous):

a = 8 a + 8d = 80 8 + 8d = 80 d = 9 a + 25d = 8 + 25(9) = 233

OpenStudy (anonymous):

@moneybird 25th term will be "a+24d" i think

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

25th term is 224

OpenStudy (lilg132):

oops

OpenStudy (lilg132):

a + 24d = 8 + 24(9) = 8 + 216 = 224 is the 25th term

OpenStudy (anonymous):

medal medal

OpenStudy (anonymous):

224

OpenStudy (ash2326):

a9=a1+(9-1)D a1=FIRST TERM D=common difference 80=8+8*D D=9 a25=a1+(25-1)D a25=8+24*9 a25=224

OpenStudy (anonymous):

add me at http://www.facebook.com/profile.php?id=100002968793465